# Complements of caustics of the real $J_{10}$ singularities

### v3

@@ 4. Proof of Proposition 9
Proposition 9 follows from this corollary in the same way as Proposition 3 of [15] was deduced from Lemmas 4 and 5 and Proposition 22. Namely, using the Lyashko–Looijenga covering, we can assume that our two Morse polynomials $f,\tilde{f}$ have equal sets of critical values. Using the diffeomorphisms of class $\mathfrak{G}$} we can also assume that they lie in the space of deformation (1) or (2). We draw a generic path in this space, connecting the polynomial $f$ with a point in the distinguished stratum $(E_{8},A_{2})$ or $(D_{8}^{-},A_{2})$ considered in Corollary 19. Using the Lyashko–Looijenga covering (and the coincidence of virtual Morse functions associated with $f$ and $\tilde{f}$), we can uniquely draw a path starting from the other polynomial $\tilde{f}$, repeating the sets of critical values of the functions of the first path and undergoing the same standard surgeries. The endpoint of the latter path can only either coincide with the endpoint of the first path, or be symmetric to it by the reflection (4). In a small neighborhood of these endpoints, the statement of Proposition 9 is trivial. By the continuity along our paths, this also holds for their starting points. $\Box$
5. Realization of some virtual Morse functions by polynomials
@@
Almost all realizations given in this section are invariant under the involution (4) and hence prove the achirality of the corresponding components.
Remark 20.
@@
In all our pictures, the $y$ axis is horizontal and $x$ axis is vertical and directed upwards.
5.1. $\Phi_{1}$
@@
in particular the corresponding component is achiral. The zero sets of these two polynomials are outlined in the left parts of Figs. 3 and 4.
Virtual Morse functions with eight real critical points and exactly two (respectively, one, none) maximum points can be realized by polynomials whose zero level sets are shown in Fig. 18 left (respectively, center, and right).
@@
(12) $x\left((y^{2}-1)^{2}+x^{2}-\varepsilon\right),\quad\varepsilon\in(0,1),$
For the second, we can take the polynomial
@@
To construct the third polynomial, we take the perturbation $x^{3}+y^{6}+\varepsilon(y^{5}+xy^{4})$, $\varepsilon>0$, of singularity $x^{3}+y^{6}$. It has one critical point of class $E_{8}$ and no other real critical points. Then we apply the standard perturbation of this $E_{8}$ singularity, as described on page 17 of [1].
@@
(14) $x^{3}+y^{6}-\varepsilon xy^{2}$
of $J_{10}^{1}$ singularity $x^{3}+y^{6}$ has a critical point of class $D_{4}^{-}$, two local minima and no other critical points, see Fig. 19 (center). Additional perturbations of its $D_{4}^{-}$ critical point, shown in Fig. 19 left and right can be performed as indicated in [1], [14] preserving the symmetry with respect to the coordinate $y$. These perturbations realize virtual Morse functions with six real critical points, exactly one (respectively, none) of which are maxima. The perturbation
@@
has only two real critical points: a minimum and a saddle.
@@
5.2. $\Phi_{3}$
Two polynomials of class $\Phi_{3}$ with ten real critical points and one maximum point, realizing the D-graphs shown in Figs. 10 and 11, can be chosen as
@@ 5.2. $\Phi_{3}$
Their zero sets are shown in the left parts of the same figures; the first of them is invariant under the involution (4). The polynomial of class $\Phi_{3}$ with ten real critical points, exactly two of which are the maxima, can be chosen as follows:
@@
A polynomial of class $\Phi_{3}$ with eight real critical points, exactly one of which is a maximum, is given by
(22) $x(x+(y-1)^{2}-2)(x+(y+1)^{2}-2)+\varepsilon y^{6}\ .$
The zero level set of this polynomial without the term $\varepsilon y^{6}$ is shown in Fig. 21.
@@
(23) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2)+\varepsilon y^{6},$
it realizes the class with Card = 131148. The left picture of Fig. 22 shows the zero set of this polynomial without the term $\varepsilon y^{6}$. The second virtual component has Card = 82350 and probably has no symmetric realizations. Nevertheless it is achiral, see § 6.1.The zero set of some polynomial representing it is given in Fig. 22 (right). A realization of this picture is as follows: first we take the perturbation
$\frac{1}{3}(x^{3}-xy^{4})-\varepsilon x^{2}y$
of the original $J_{10}^{3}$ singularity $\frac{1}{3}(x^{3}-xy^{4})$, and then perturb its single real critical point of type $D_{8}^{-}$ as shown on p. 15 of [1].
An isotopy invariant proving that these two Morse functions indeed belong to different isotopy classes is as follows. For any function $f$ of class $\Phi_{3}$ with exactly three local minima, a unique parabola (or line) of the form $x=ay^{2}+by+c$ exists that passes through these three points. The restriction of $f$ to this curve is a polynomial of degree at most six in the coordinate $y$. Since it has three minima, it must be of degree exactly six. Therefore, the coefficient $a$ is not a root of the polynomial $\alpha t^{3}+\beta t^{2}+\gamma t+\delta$ where $\alpha x^{3}+\beta x^{2}y^{2}+\gamma xy^{4}+\delta y^{6}$ is the principal quasihomogeneous part of $f$. Thus, the position of $a$ among three roots of this polynomial is an invariant of the isotopy class. For the left picture of Fig. 22 the coefficient $a$ does not separate these roots, while for the right picture it separates one root from the other two.
@@
The isotopy components of class $\Phi_{3}$ with exactly six real critical points, exactly one or none of which is a local maximum, are represented by the polynomials
@@
with sufficiently small $\varepsilon>0$. Their zero-level sets are shown in Fig. 23. The degeneration of the second of these polynomials at $\varepsilon=0$ has only two real critical points of type $A_{3}$.
@@
A polynomial of class $\Phi_{3}$ with exactly four real critical points and no local maxima is given by
@@
Consider a virtual component on the formal graph of class $\Phi_{1}$ or $\Phi_{3}$.
Denote by $\mathfrak{H}$} the one-dimensional simplicial cochain of this virtual component with coefficients in ${\mathbb{Z}}_{2}$, which takes the non-zero value on the edges corresponding to the elementary operation $s2$ of changing the order of two neighboring real critical values, both of which are achieved at minimum points.
@@
If all virtual Morse functions of a virtual component have exactly two or three minima, then this component is chiral if and only if the cocycle $\mathfrak{H}$} is trivial on it.
Proof. Let $N=2$ or $3$ be this number of minima of all these functions. Consider a generic path in the space of Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$ that connects two mutually symmetric polynomials $f(x,y)$ and f(x,y)f(x,-y). The endpoints of this path are associated with the same virtual Morse function; therefore, the associated path in the formal graph is a cycle. Polynomials of class $\Phi_{1}$ or $\Phi_{3}$ cannot have two minima on the same vertical line. Therefore, our path preserves the order of the projections of the minimum points to the $y$ axis. On the other hand, for any $i\in[1,N]$ the critical value of the $i$th in this order minimum point of the starting polynomial of this path is equal to the critical value of the (Ni)(N-i)-th minimum point of the final polynomial of the path. Thus, our cycle contains an odd number of $s2$ surgeries that change the order of the critical values at the minimum points.
Conversely, given a cycle in a virtual component of the formal graph, let us realize it by a path in the space $\Phi_{1}$ or $\Phi_{3}$ of real polynomials starting from an arbitrary vertex of the cycle. That is, the path should start with a polynomial associated with this vertex and consecutively experience all surgeries encoded by the edges of the cycle. The existence of such a path is ensured by Proposition 7. If the cocycle $\mathfrak{H}$} takes the non-zero value on the initial cycle, then the isotopy classes of the endpoints of this realized path are different. $\Box$
@@
We could formulate the same criterion in terms of the critical values at the maximum points. Therefore, the other cocycle defined in this virtual component as $\mathfrak{H}$} but in the terms of the maxima instead of the minima is homologous to $\mathfrak{H}$}.
Almost all virtual components for which Theorem 15 claims their achirality are represented by polynomials that are symmetric with respect to the involution (4), see (11), (12), (13), (14), (15), (16), (17), (18), (19), (22), (23), (24), (25), (26), and (27). The remaining two virtual Morse functions are of class $\Phi_{3}$ with eight real critical points and Card = 82350; one of them has no maxima (see Fig. 22 right) and the other is formed from it by the virtual analog of involution (5). The simplicial cocycle $\mathfrak{H}$} takes the non-zero value on the cycle $A\to B\to D\to C\to A$ in the corresponding virtual component, see Fig. 2 and Example 5. Indeed, this cocycle takes the non-zero value only on the edge $[A,C]$ of this cycle. This proves the achirality of these two virtual components as well.
@@
This cycle was found by a small suspension of our program.

### v5

@@ 4. Proof of Proposition 9
Using the Lyashko–Looijenga covering, we can assume that our two Morse polynomials $f$ and $\tilde{f}$ have equal sets of critical values and equal systems of paths in ${\mathbb{C}}^{1}$ defining the vanishing cycles. Using the diffeomorphisms of class $\mathfrak{G}$} we can also assume that they lie in the space of deformation (1). We draw a generic piecewise-algebraic path $I:[0,1]\to\Phi_{1}$ in this space, connecting the polynomial $f\equiv I(0)$ with one of two points of the $(E_{8},A_{2})$ stratum considered in Corollary 20. Using the Lyashko–Looijenga map and the coincidence of virtual Morse functions associated with $f$ and $\tilde{f}$, we can uniquely draw a path $\tilde{I}:[0,1]\to\Phi_{1}$ starting from the polynomial $\tilde{f}$, repeating the sets of critical values of the functions of the first path and undergoing the same standard surgeries. The endpoint $\tilde{I}(1)$ of this path has the critical points of classes $E_{8}$ and $A_{2}$ with the same critical values as $I(1)$, so according to Corollary 20 it either coincides with the point $I(1)$ or is symmetric to it via the involution (4).
In the second case, when $I(1)\neq\tilde{I}(1)$, we apply this involution to entire path $\tilde{I}$, including its starting point $\tilde{f}$, and reduce the proof to the first case.
In the first case, the final segments of the paths $I$ and $\tilde{I}$ should coincide in a neighborhood of their endpoint $I(1)=\tilde{I}(1)$. Indeed, suppose that the points $I(1-\varepsilon)$ and $\tilde{I}(1-\varepsilon)$ are different for small $\varepsilon>0$. We can assume that these points lie in the set of generic polynomials. Consider the map from one of these points to the other. Since their associated virtual Morse functions and critical values are the same, the covering homotopy of the Lyashko–Looijenga map extends this maps to an automorphism of the space $\Phi_{1}$, sending the point $I(1)$ to itself and commuting with the Lyashko–Looijenga map.
Analyzing the multisingularity $(E_{8},A_{2})$ as in § 3.5 of [20] proves that such an automorphism can only be the identity. Indeed, by the versality property, a neighborhood of the point $I(1)$ is naturally diffeomorphic to the Cartesian product of the parameter spaces of the canonical versal deformations of the simple singularities of classes $E_{8}$ and $A_{2}$. Namely, in this neighborhood the family (1) can be considered as a deformation of the $E_{8}$ singularity of the polynomial $I(1)$, and also as a deformation of $A_{2}$ singularity of the polynomial $I(1)+\frac{1}{5}$. The projection maps of the Cartesian product structure to its factors are maps of the parameter spaces that locally induce the deformation (1) from these two deformations as in the definition of versal deformations, see § 8 in [4]. The aforementioned automorphism of a neighborhood of the point $I(1)$ then induces the automorphisms of the canonical versal deformations of the $E_{8}$ and $A_{2}$ singularities that commute with the corresponding Lyashko–Looijenga coverings. All complex automorphisms of versal deformations of simple singularities that commute with the Lyashko–Looijenga covering are listed in [9]. This list implies that there are no such non-trivial automorphisms that preserve the space of real functions. Thus, the paths $I$ and $\tilde{I}$ coincide near their common endpoint, and hence they coincide everywhere, including their starting points $f$ and $\tilde{f}$. This proves Proposition 9 for polynomials of class $\Phi_{1}$.
In the case $\Phi_{3}$, an additional difficulty occurs at the final part of the proof. Namely, the canonical real versal deformation
(11) $F(\xi,\lambda)\equiv\xi^{2}\eta-\eta^{7}+\lambda_{1}+\lambda_{2}\xi+\lambda_{3}\eta+\lambda_{4}\eta^{2}+\lambda_{5}\eta^{3}+\lambda_{6}\eta^{4}+\lambda_{7}\eta^{5}+\lambda_{8}\eta^{6}$
of $D_{8}^{-}$ singularity has exactly one non-trivial automorphism commuting with the Lyashko–Looijenga map: it is defined by the reflection
(12) $(\xi,\eta)\leftrightarrow(-\xi,\eta).$
This fact follows from the description of all complex automorphisms of the $D_{8}$ singularities that commute with the Lyashko–Looijenga covering (see [9]), and from the selection of automorphisms that preserve the set of real polynomials.
Therefore, a neighborhood of the point $I(1)$ in the space of polynomials (2) has exactly one such automorphism as well. Specifically, this neighborhood can be naturally identified with the direct product of the parameter spaces of the standard versal deformations of the $D_{8}^{-}$ and $A_{2}$ singularities. The unique nontrivial automorphism of the neighborhood of the point $I(1)$ acts as the product of the automorphism on the eight-dimensional factor induced from by the reflection (12) and the identity automorphism of the versal deformation of the $A_{2}$ singularity.
Let us prove that this unique nontrivial local automorphism does not preserve the associated virtual Morse functions.
Denote by $U$ the set of polynomials of the form (2) near the point $I(1)$, such that the $D_{8}^{-}$-point of the polynomial $I(1)$ splits into three minima with negative critical values and five saddlepoints with positive critical values, and the $A_{2}$-point splits into two real critical points, a minimum and a saddlepoints with critical values near $\frac{1}{5}$. According to Theorem 1.7 of [11] and the direct product structure described above, this set consists of a single contractible connected component. We can assume that the path $I$ approaches the point $I(1)$ from inside this domain $U$.
---678910
---678109
Figure 19. Ordering the critical values of perturbations of $D_{8}^{-}$ singularity
The boundary of the domain $U$ contains a Morse polynomial considered in Lemma 22. Therefore, the D-graphs of all polynomials from $U$ also are as shown in Fig. 15 (right).
The perturbations of $I(1)$ that satisfy the conditions of Lemma 22 are not generic, as their five Morse points have common critical value 0. Such a perturbation $f_{\lambda}$ can be slightly perturbed again so that the resulting polynomial $f_{\lambda^{\prime}}$ is generic and belongs to the set $U$, and the critical values at all its five saddlepoints obtained by splitting the $D_{8}^{-}$ singularity are ordered in ${\mathbb{R}}^{1}$ as shown by the numbers in Fig. 19 (left) at the corresponding saddlepoints of the polynomial $f_{\lambda}$. (These numbers start with 6 since the lower critical values at the other five critical points are taken into account.) Let $f_{\lambda^{\prime\prime}}$ be the Morse polynomial obtained from $f_{\lambda^{\prime}}$ by our local automorphism. It is very close to $f_{\lambda}$ and hence to $f_{\lambda^{\prime}}$. The critical values of $f_{\lambda^{\prime\prime}}$ are ordered as shown in Fig. 19 (right). According to the Gusein-Zade–A’Campo calculation method of intersection indices, the vanishing cycles in the sets $f^{-1}_{\lambda^{\prime}}(0)$ and $f^{-1}_{\lambda^{\prime\prime}}(0)$ that vanish at the critical points labeled by 9 and 10 correspond to the horns of the Coxeter–Dynkin graph 678109 of the $D_{8}^{-}$ singularity.
Two generic perturbations $f_{\lambda^{\prime}}$ and $f_{\lambda^{\prime\prime}}$ can be continuously deformed into each other within the domain $U$ in such a way that the cycle vanishing at the point marked by 9 in Fig. 19 (left) will be moved to the cycle vanishing at the point marked by 10 in Fig. 19 (right), and the orders of critical values of all other critical points of $f_{\lambda}$ will not change. Therefore, if the perturbations $f_{\lambda^{\prime}}$ and $f_{\lambda^{\prime\prime}}$ have equal associated virtual Morse functions, then the intersection indices of both vanishing cycles $\Delta_{9}$ and $\Delta_{10}$ in $f^{-1}_{\lambda^{\prime}}(0)$ with all other basic vanishing cycles should be the same. However, Fig. 15 (right) shows that the cycle corresponding to one horn of the $D_{8}$-subgraph of the D-graph has nonzero intersection indices with both cycles corresponding to the vertices of the $A_{2}$-subgraph, while the cycle corresponding to the other horn has nonzero intersection index with only one of them. Thus, the virtual Morse functions associated with polynomials $f_{\lambda^{\prime}}$ and $f_{\lambda^{\prime\prime}}$ are different. Therefore, the paths $I$ and $\tilde{I}$ coincide near their endpoints and, consequently, everywhere, including their initial points $f$ and $\tilde{f}$. $\Box$
5. Realization of virtual Morse functions
@@
In this section, we realize by polynomials many virtual components, whose existence was predicted in Theorems 14 and 15. Almost all these realizations are invariant under the involution (4) and hence prove the achirality of the corresponding isotopy classes.
@@
In all our pictures, it is assumed that the $y$ axis is horizontal and the $x$ axis is vertical and directed upwards.
5.1. Realizations of virtual Morse functions of class $\Phi_{1}$
@@
in particular the corresponding connected component of the space of Morse polynomials is achiral. The zero-level sets of these two polynomials are outlined in the left parts of Figs. 3 and 4.
Remark 23.
The polynomials (13) and (14) are not generic since they have multiple critical value 0. However, their arbitrary small generic perturbations realize the promised virtual components, cf. Definition 8.
Virtual Morse functions with eight real critical points and exactly two (respectively, one, respectively, no) maximum points can be realized by polynomials whose zero-level sets are shown in Fig. 20 left (respectively, Fig. 20 right, respectively, Fig. 18 left).
@@
(15) $x\left(x^{2}+(y^{2}-1)^{2}-\varepsilon\right),\quad\varepsilon\in(0,1),$
For the second, we take the polynomial
@@
To construct the third polynomial, we take the perturbation
(17) $x^{3}+y^{6}+\varepsilon(x^{2}y+3y^{5}),\quad\varepsilon>0,$
of singularity $x^{3}+y^{6}$. It has one critical point of class $E_{8}$ and no other real critical points. Then we apply the standard perturbation of this $E_{8}$ singularity, as described on page 17 of [1]. This perturbation can be chosen arbitrarily small, so that it does not return the non-real critical points of (17) to the real domain.
@@
(18) $x^{3}+y^{6}-\varepsilon xy^{2},\quad\varepsilon>0,$
of $J_{10}^{1}$ singularity $x^{3}+y^{6}$ has a critical point of class $D_{4}^{-}$, two local minima and no other critical points, see Fig. 21 (center). Additional perturbations of its $D_{4}^{-}$ critical point, shown in Fig. 21 (left and right), can be performed as indicated in [1] or [19] preserving the symmetry with respect to the coordinate $y$. These perturbations realize virtual Morse functions with six real critical points, exactly one (respectively, none) of which are maxima. The perturbation
@@
has only two real critical points: a minimum and a saddlepoint, see Fig. 22 (right).
@@
5.2. Realizations of virtual Morse functions of class $\Phi_{3}$
Two polynomials of type $\Phi_{3}$ with ten real critical points, exactly one of which is a maximum point, can be chosen as
@@ 5.2. Realizations of virtual Morse functions of class $\Phi_{3}$
They realize the D-graphs shown in Figs. 10 and 11. Their zero-level sets are shown in the left parts of the same figures. The first of these polynomials is invariant under the involution (4).
A polynomial of type $\Phi_{3}$ with ten real critical points, exactly two of which are the maxima, can be chosen as follows:
@@
A polynomial of type $\Phi_{3}$ with eight real critical points, exactly one of which is a maximum, is given by
(26) $x(x+(y-1)^{2}-2)(x+(y+1)^{2}-2)+\varepsilon y^{6}\equiv x^{3}+2x^{2}y^{2}-2x^{2}+xy^{4}-6xy^{2}+x+\varepsilon y^{6}\$
with any sufficiently small $\varepsilon>0$. The zero level set of this polynomial without the last term $\varepsilon y^{6}$ (which makes its principal quasihomogeneous part non-degenerate of class $J_{10}^{3}$) is shown in Fig. 23.
@@
(27) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2)+\varepsilon y^{6}\equiv x^{3}+2x^{2}y^{2}+4x^{2}+xy^{4}-12xy^{2}+4x+\varepsilon y^{6}.$
It realizes the class with Card = 131148. The Fig. 24 (left) shows the zero-level set of this polynomial without the term $\varepsilon y^{6}$. The second isotopy class has Card = 82350 and no symmetric realizations. Nevertheless, it is achiral, see § 6.1. A realization of this class is shown in Fig. 24 (right). To construct it, we first take the perturbation
$x^{3}-xy^{4}-\varepsilon x^{2}y$
of the original $J_{10}^{3}$ singularity $x^{3}-xy^{4}$, and then perturb its single real critical point of type $D_{8}^{-}$ as shown on p. 15 of [1].
An isotopy invariant proving that these two Morse functions indeed belong to different isotopy classes is as follows. For any polynomial $f$ of type $\Phi_{3}$ with exactly three local minima, a unique parabola (or line) with equation of the form $x=ay^{2}+by+c$ exists that passes through these three points. The restriction of $f$ to this curve is a polynomial of degree at most six in the coordinate $y$. Since it has three minima, it is of degree exactly six. Therefore, the coefficient $a$ of this parabola is not a root of the polynomial $\alpha t^{3}+\beta t^{2}+\gamma t+\delta,$ where $\alpha x^{3}+\beta x^{2}y^{2}+\gamma xy^{4}+\delta y^{6}$ is the principal quasihomogeneous part of $f$. Thus, the position of the number $a$ among three roots of this polynomial is an invariant of the isotopy class. For a polynomial realizing the left picture of Fig. 24 the coefficient $a$ does not separate these roots, while for the right picture it separates one root from the other two.
@@
The isotopy component of type $\Phi_{3}$ with exactly six real critical points, exactly one (respectively, none) of which is a local maximum, is represented by the polynomial
@@
with sufficiently small $\varepsilon>0$). The zero-level sets of these polynomials are shown in Fig. 25. The degenerate version of the second of these polynomials corresponding to $\varepsilon=0$ has only two real critical points of type $A_{3}$.
@@
A polynomial of type $\Phi_{3}$ with exactly four real critical points and no local maxima is given by
@@
Denote by $\mathfrak{H}$} the one-dimensional simplicial cochain of the formal graph of type $\Phi_{1}$ or $\Phi_{3}$ with coefficients in ${\mathbb{Z}}_{2}$, which takes the non-zero value only on the edges corresponding to the type $s2$ elementary operation of changing the order of two neighboring real critical values, both of which are achieved at minimum points.
@@
If all virtual Morse functions in a virtual component of $\Phi_{1}$ or $\Phi_{3}$ type have exactly two or three minima, then this component is chiral if and only if the cocycle $\mathfrak{H}$} is trivial on it.
Proof. Let $N$ be the number of minima of these virtual functions. Suppose that the virtual component is achiral and hence is associated with only one isotopy class of Morse polynomials in $\Phi_{1}$ or $\Phi_{3}$. Then, there exists a generic path within this isotopy class that connects two mutually symmetric generic polynomials, $f(x,y)$ and f(x,y)f(x,-y). The endpoints of this path are associated with the same virtual Morse function. Therefore, the associated path in the formal graph is a cycle. Polynomials of classes $\Phi_{1}$ and $\Phi_{3}$ cannot have two minima on the same vertical line $\{x=c\}$ because they are of degree three in the $x$ variable. Thus, our path preserves the order of the projections of the minimum points to the $y$ axis. Consider the following two permutations of the numbers $1,2,\dots,N$. The permutation $\sigma$ depends only on the initial polynomial $f$ and maps each number $i$ to the order of the critical value of $f$ at the $i$-th minimum point from the left among all such critical values. The permutation $\varkappa$ is the permutation of orders of critical values at minima defined by the continuation along our path. The elementary transpositions that form the permutation $\varkappa$ along this path correspond to the edges of the associated path in the virtual component, on which the cochain $\mathfrak{H}$} takes the non-zero value.
For any $i\in[1,N]$ the critical value of the $i$th minimum point from the left of the initial polynomial of this path is equal to the critical value of the (Ni)(N-i)-th minimum point of the final polynomial. Therefore, the permutations $\varkappa$ and $\left(\begin{array}[]{ccccc}1&2&\dots&N-1&N\\ N&N-1&\dots&2&1\end{array}\right)$ are conjugate via $\sigma$, in particular, have the same parity. For $N=2$ or 3, this parity is odd, and hence the cocycle $\mathfrak{H}$} takes a non-zero value on the cycle.
Conversely, given a cycle in a virtual component of the formal graph, we can realize it by a path in the space $\Phi_{1}$ or $\Phi_{3}$ of real polynomials starting from an arbitrary vertex of the cycle. That is, the path begins with a polynomial associated with this vertex and successively undergoes all surgeries encoded by the edges of the cycle. The existence of such a path is ensured by Proposition 7. If the cocycle $\mathfrak{H}$} takes the non-zero value on the initial cycle, then the isotopy classes of the endpoints of the realized path in the space of Morse polynomials differ because a closed loop in $\Phi_{1}$ or $\Phi_{3}$ defines the trivial permutation $\varkappa$. $\Box$
@@
We could formulate the same criterion in terms of the critical values at the maximum points. Therefore, if both inertia indices $m_{-}$ and $m_{+}$ of the passport invariant of the virtual component are equal to two or three, then another cocycle defined in this virtual component, similar to $\mathfrak{H}$} but in the terms of maxima instead of minima, is homologous to $\mathfrak{H}$}.
Almost all of the virtual components that Theorem 16 claims are achiral are represented by symmetric polynomials with respect to the involution (4), see the polynomials (14)–(16), (18)–(23), and (26)–(31). The remaining two virtual Morse functions are of type $\Phi_{3}$ with eight real critical points and Card = 82350. One of them has no maxima (see Fig. 24 right) and the other is obtained from it by the involution (5). The simplicial cocycle $\mathfrak{H}$} takes the non-zero value on the cycle $A\to B\to D\to C\to A$ in the corresponding virtual component, see Fig. 2 and Example 5. Indeed, this cocycle takes the non-zero value only on the edge $[A,C]$ of this cycle. This proves the achirality of these two virtual components as well.
@@
This cycle was found by a minor suspension of our program.
Remark 27.
Another virtual component with eight real critical points, no maxima and Card=131148 contains the virtual Morse function
[…]
In this virtual component, a cycle that proves its achirality consists of only $s2$ type edges. It defines the transpositions of critical values at the critical points 1 and 3, 5 and 6, and 4 and 7. See also Fig. 24 (left).